# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

### v3

@@ B.2 Training Hyperparameters
B.3 Loss Functions
For forecasting tasks, models were trained using quantile loss to produce probabilistic forecasts across multiple quantile levels. For scientific parameter inference tasks (e.g., estimating $R_{0}$), quantile loss was also used to generate full predictive intervals rather than point estimates. For classification tasks (source identification), cross-entropy loss was used. No auxiliary losses were applied in any experiments.
@@ Appendix D Back-to-Simulation Attribution
To enable mechanistically interpretable diagnostics, SGNNs support back-to-simulation attribution, a retrieval-based method that identifies the most similar synthetic regimes to a given real-world input based on the SGNN’s internal representation space. This section describes the implementation in full detail.
@@
To perform attribution, a real-world time series (e.g., weekly COVID-19 deaths in Michigan) is first transformed into the format expected by the pretrained SGNN encoder. This involves applying the exact same preprocessing used during model training—for example, in our case, log1p transformation followed by z-score normalization using training-set statistics. The model-specific context length is enforced, with earlier time points padded if necessary. The processed input is then passed through the full SGNN encoder, which, in this case, includes multi-scale convolutional feature extraction, categorical embeddings (e.g., disease type, time-series type, calendar indices), and a CNN-Transformer backbone. The encoder output is pooled using a validity mask over the time dimension to yield a fixed-length embedding $z\in\mathbb{R}^{d}$ (with $d=1024$ in our primary implementation). While preprocessing steps and model architecture can vary by domain, attribution requires that embeddings be computed using the same pipeline and parameters as used during training.
@@ Appendix E Additional Results
E.1 Ecological Forecasts
To illustrate the forecasting capabilities of SGNNs in ecological systems, we provide a visual comparison of model predictions against ground-truth population dynamics in the lynx-hare predator-prey system. Figure 5 shows forecast trajectories from SGNNs, Rosenzweig-MacArthur growth model (RMG), a purely mechanstic model, and vector autoregressive moving average models (VARMA), a purely data-driven model, across four rolling prediction windows.
These results emphasize the capability of simulation-grounded pretraining for capturing complex ecological feedbacks and suggest that SGNNs are capable of robust extrapolation even in systems with limited historical data and strong cyclical behavior.
Figure 5: SGNNs outperform classical models in forecasting lynx-hare predator-prey dynamics. Rolling forecasts are shown for four evaluation windows, comparing SGNN (red) to mechanistic models (RMG, blue) and VARMA (green) models. True population trajectories are plotted in black. SGNNs consistently maintain accurate phase and amplitude tracking, while baselines deteriorate.
E.2 State $R_{0}$ Estimates
@@
  • General interpretability: SGNN’s estimates tend to reflect known drivers of early outbreak heterogeneity—population density, travel exposure, and reporting infrastructure—despite having no access to these covariates. This suggests that SGNN’s latent dynamics have internalized meaningful global patterns purely from time-series shape and signal-to-noise characteristics.
Overall, these results demonstrate SGNN’s ability to produce informative and accurate $R_{0}$ estimates directly from early outbreak curves. This capability—learning to infer latent, unobservable scientific quantities with calibrated uncertainty—would be impossible using traditional supervised learning. It highlights the potential for simulation-grounded models to support real-time scientific inference during future emergent crises.

### v4

@@ B.2 Training Hyperparameters
B.3 Loss Functions and Optimization Objectives
For each task, we used loss functions tailored to the specific requirements. For forecasting and regression, these objectives are designed to produce not only point estimates but also estimate the underlying uncertainty.
Probabilistic Forecasting and Parameter Inference.
For time-series forecasting and the estimation of unobservable latent parameters (e.g., $R_{0}$), we used the quantile loss (also often called pinball loss). This objective allows the SGNN to produce a predictive distribution rather than a single value. For a target $y$ and a predicted quantile $\hat{y}_{\tau}$ at probability level $\tau\in(0,1)$, the loss is defined as:
$L_{\tau}(y,\hat{y}_{\tau})=\max(\tau(y-\hat{y}_{\tau}),(1-\tau)(\hat{y}_{\tau}-y))$ (7)
To generate calibrated 95% prediction intervals, we optimize the model over a discrete set of $Q$ quantiles (e.g., $\tau\in\{0.025,0.5,0.975\}$). The total objective is the mean across all quantiles:
$\mathcal{L}_{\text{forecast}}=\frac{1}{Q}\sum_{q=1}^{Q}L_{\tau_{q}}(y,\hat{y}_{\tau_{q}})$ (8)
Source Identification (Classification).
We treat the identification of diffusion sources as a multi-class classification problem over $N$ nodes in the graph. The Graph Attention Network (GAT) outputs a probability distribution $\mathbf{\hat{y}}$ via a softmax activation. We minimize the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}y_{i}\log(\hat{y}_{i})$ (9)
where $\mathbf{y}$ is the one-hot encoded ground-truth source vector. This objective encourages the model to assign maximum likelihood to the true mechanistic origin of the observed cascade.
Chemical Yield Regression (Multi-Head).
For the Suzuki-Miyaura yield prediction, we employ a composite objective that balances the predictions of the hybrid wide-and-deep architecture with an uncertainty estimation head. The total loss $\mathcal{L}_{\text{chem}}$ is defined as:
$\mathcal{L}_{\text{chem}}=\|y-\hat{y}_{\text{ens}}\|_{2}^{2}+\alpha\|y-\hat{y}_{a}\|_{2}^{2}+\beta\frac{1}{N}\sum_{i=1}^{N}\hat{c}_{i}|y_{i}-\hat{y}_{\text{ens},i}|$ (10)
where $\hat{y}_{\text{ens}}=\sum_{k\in\{m,a\}}w_{k}\hat{y}_{k}$ is the weighted ensemble of the main ($m$) and auxiliary ($a$) heads, and $w_{k}$ are softmax-normalized learnable weights. The second term, scaled by $\alpha$, ensures predictive utility in the auxiliary head to promote representation diversity. The final term, scaled by $\beta$, incorporates the confidence head output $\hat{c}\in[0,1]$, which acts as a learned weight on the absolute residuals to capture heteroscedastic noise and failure-prone regimes.
@@ Appendix D Back-to-Simulation Attribution
To enable mechanistically interpretable diagnostics, SGNNs support back-to-simulation attribution, a retrieval-based method that identifies the most similar synthetic regimes to a given real-world input based on the SGNN’s internal representation space. This section describes the implementation in more detail.
@@
To perform attribution, a real-world time series (e.g., weekly COVID-19 deaths in Michigan) is first transformed into the format expected by the pretrained SGNN encoder. This involves applying the exact same preprocessing used during model training. For example, in our case, log1p transformation followed by z-score normalization using training-set statistics. The model-specific context length is enforced, with earlier time points padded if necessary. The processed input is then passed through the full SGNN encoder, which, in this case, includes multi-scale convolutional feature extraction, categorical embeddings (e.g., disease type, time-series type, calendar indices), and a CNN-Transformer backbone. The encoder output is pooled using a validity mask over the time dimension to yield a fixed-length embedding $z\in\mathbb{R}^{d}$ (with $d=1024$ in our primary implementation). While preprocessing steps and model architecture can vary by domain, attribution requires that embeddings be computed using the same pipeline and parameters as used during training.
@@ Appendix E Additional Results
E.1 Infectious Disease Forecast Evaluation
To complement the forecasting skill analysis presented in the main text, we evaluated all models using standard probabilistic scoring metrics from infectious disease forecasting: the relative Weighted Interval Score (WIS) [5], 50% prediction interval coverage, and 95% prediction interval coverage.
The Weighted Interval Score is a proper scoring rule that jointly evaluates the sharpness and calibration of probabilistic forecasts by averaging over a set of central prediction intervals [5]. The WIS generalizes the absolute error to the probabilistic setting: it penalizes both wide intervals (poor sharpness) and observations falling outside predicted intervals (poor calibration), and reduces to the MAE when only a point forecast is provided.
To enable comparison across locations and time periods with different scales, we report relative WIS, computed as the ratio of each model’s mean WIS to that of a persistence baseline. The persistence baseline forecasts the most recent observed value as its point prediction at all horizons, with prediction intervals derived from the historical variance of week-over-week changes in the target series. A relative WIS below 1.0 indicates improvement over the baseline, while values above 1.0 indicate worse performance. We additionally report empirical coverage of the 50% and 95% prediction intervals, where well-calibrated models should achieve coverage rates near these levels.
Results are presented in Table 2 for COVID-19 mortality forecasting and Table 3 for dengue case forecasting. On COVID-19, SGNNs matched the relative WIS of the best CDC Forecast Hub model (0.63) while achieving better 50% coverage (0.51 vs. 0.47). The PINN-based EINN model performed worse than the persistence baseline (relative WIS = 1.10) with poor calibration at both interval levels. On dengue, SGNNs again outperformed all baselines in relative WIS (0.80), with well-calibrated 95% coverage (0.90). The ETS baseline achieved comparable 95% coverage but with wider intervals reflected in its higher relative WIS (1.04).
Model Relative WIS $\downarrow$ 50% Coverage 95% Coverage
SGNN 0.63 0.51 0.91
CDC Forecast Hub (Best) 0.63 0.47 0.93
Chronos 0.88 0.42 0.82
EINN (PINN) 1.10 0.38 0.71
Table 2: Probabilistic forecast evaluation for COVID-19 mortality. Relative WIS is computed against a persistence baseline (values <1.0<1.0 indicate improvement). Nominal coverage targets are 0.50 and 0.95. Bold indicates best performance per column.
Model Relative WIS $\downarrow$ 50% Coverage 95% Coverage
SGNN 0.80 0.47 0.90
ETS 1.04 0.45 0.90
EINN (PINN) 2.03 0.31 0.64
Table 3: Probabilistic forecast evaluation for dengue case forecasting. SGNNs were evaluated fully zero-shot with no exposure to dengue data or vector-borne transmission dynamics. Relative WIS is computed against a persistence baseline.
E.2 Ecological Forecasts
To illustrate the forecasts of SGNNs in ecological systems, we provide a visual comparison of model predictions against ground-truth population dynamics in the lynx-hare predator-prey system. Figure 5 shows forecast trajectories from SGNNs, Rosenzweig-MacArthur growth model (RMG), a purely mechanstic model, and vector autoregressive moving average models (VARMA), a purely data-driven model, across four rolling prediction windows.
Figure 5: SGNNs outperform classical models in forecasting lynx-hare predator-prey dynamics. Rolling forecasts are shown for four evaluation windows, comparing SGNN (red) to mechanistic models (RMG, blue) and VARMA (green) models. True population trajectories are plotted in black. SGNNs maintain more accurate phase and amplitude tracking versus baselines.
E.3 State $R_{0}$ Estimates
@@
  • General interpretability: SGNN’s estimates tend to reflect known drivers of early outbreak heterogeneity (population density, travel exposure, and reporting infrastructure) despite having no access to these covariates, suggesting that the SGNN can capture these patterns just from time-series shape and signal-to-noise characteristics.